# Demystifying Language Model Forgetting with Low-rank Example Associations

### v3

@@ 3.2 What are More Fine-Grained Associations?
Figure 4: Patterns in associations captured by $k$-th component in SVD of the assoication matrix $Z$, exemplified with OLMo-7B-Instruct (full-parameter tuning, also in Fig. 2(c)). The highlighted regions display patterns that TruthfulQA tasks cause more forgetting ($k=2$) and MMLU causes more forgetting on certain tasks in FLAN v2 and Open-Orca ($k=3$).
We extract more fine-grained associations where fine-tuning on a task causes a specific set of upstream examples to be forgotten. We perform Singular Value Decomposition (SVD) over the association matrices $Z$ and progressively reconstruct $Z$ with up to $r$-th singular values and vectors as $Z_{r}=\sum_{k=1}^{r}s_{k}\bm{\alpha}_{k}\bm{\beta}_{k}^{T}$. $Z_{r}$ is also the optimal rank-$r$ matrix that minimizes the Frobenius norm $||Z-Z_{r}||_{F}$ and the $R^{2}$ score according to the property of SVD. The first component $Z_{1}$ corresponds to the multiplicative model we examined in Sec. 3.1; the rest of components $Z_{2..M}$ captures fine-grained associations in $Z$.
SVD reveals fine-grained associations. We present $R^{2}$ of fitting the assoication with $Z_{r}$ with progressively larger rank $r$ in Figure 3 (b). We notice that for OLMo-7B and OLMo-7B-Instruct (full-parameter fine-tuning), $R^{2}$ quickly increases to 0.69 and 0.78 with $Z_{3}$ (reconstruction with up to the third component). We visualize the matrices reconstructed from each component $s_{k}\bm{\alpha}_{k}\bm{\beta}_{k}^{T}$ in SVD of $Z$. Figure 4 provides an example of patterns captured by the $k$-th component in OLMo-7B-Instruct. We demonstrate how the visualization extracts patterns of forgetting that is conditional on the learned tasks. In Appendix D, we further examine interpretable patterns in the decomposition of $Z$ in OLMo-1B and 7B.
3.3 How do the Similarity Measures Interpret the Associations?
@@
We examine whether the associations between learned tasks and forgotten examples are interpretable from the similarity between the learned tasks and upstream examples. We consider (1) heuristic similarity measures, such as textual similarity, and (2) theoretically grounded approximations, such as inner products of gradients.
Inner products between projected gradients and model weight updates. The increase of the log perplexity ${z_{ij}}$ (also known as the cross-entropy loss) can be approximated with inner products $z^{\textrm{g-w}}_{ij}=\langle\nabla_{\theta}f(x_{j}),\theta_{T_{i}}-\theta_{0}\rangle$ under first-order Taylor expansion (Lee et al., 2019; Doan et al., 2020), where $\nabla_{\theta}f(x_{j})$ is the gradient of the loss of $x_{j}$ at the initial model before fine-tuning, and $\theta_{T_{i}}-\theta_{0}$ are the updates in the model weights after fine-tuning. Following Park et al. (2023); Xia et al. (2024), we use a random projection matrix $\bm{P}\sim\mathcal{N}_{|\theta|\times d}(0,1)$ to reduce the dimension of the gradients or the weight changes to save the cost of storing pre-computed statistics, which preserves the inner products with high probability (Johnson & Lindenstrauss, 1984).
@@
Table 1: Correlations between various measures of similarity and the actual forgetting on upstream examples after fine-tuning on a task.
@@
Gradient-Weight Diff. $z^{\textrm{g-w}}$0.003 -0.009
Gradient-Gradient Diff. $z^{\textrm{g-g}}$ 0.061 0.052
Forgetting correlates poorly with similarity measures of learned tasks and upstream examples. We evaluate correlations between the actual forgetting $z_{ij}$ and the various similarity measures $\{z^{\textrm{text}}_{ij},z^{\textrm{g-w}}_{ij},z^{\textrm{g-g}}_{ij}\}$ on OLMo-1B and summarize the results in Table $1$The table is updated from the earlier preprint (https://arxiv.org/abs/2406.14026v2) due to an implementation issue.. We notice that none of the similarity measures correlates with the actual forgetting, with a correlation $|\rho|<0.1$. An interpretation of the low correlation is that the model weights deviate from the initial weights after fine-tuning to a region where first-order approximation of forgetting does not hold. We further visualize the matrices of $z^{\textrm{g-w}}_{ij}$ and $z^{\textrm{g-g}}_{ij}$ in Fig. 8 in Appendix and provide a side-by-side comparison with the matrices of forgetting $Z$. The visualization displays distinct patterns among the three matrices. These results imply that although the association matrices $Z$ display a simple pattern, they are not well-interpreted with common similarity measures of learned tasks and forgotten examples.
@@ 4 Predicting Example Forgetting with Association Matrix Completion
We utilize our findings in Sec. 3 to predict example forgetting as the model learns a new task, a problem also studied in prior works (Jin & Ren, 2024). Finding out the most forgotten examples allows better spot of the behavior changes of the models, enabling efficient and targeted approaches to mitigate forgetting, e.g., by replaying these examples (Aljundi et al., 2019a; Wang et al., 2024). Although the ground truth forgetting can be directly obtained by running inference with the fine-tuned model over the upstream data, this requires extensive computational resources. We restrict the prediction methods to be computationally efficient.
Our analysis in Sec. 3 suggests that (1) the associations between learned tasks and forgotten examples display simple statistical patterns, while (2) the associations correlate poorly with many similarity metrics. Therefore, we hypothesize that leveraging the statistics of forgetting allows more effective prediction of forgetting than leveraging the contents of the tasks and examples. Following this intuition, we formulate prediction of example forgetting as a matrix completion problem over the empirical associations $Z$, analogical to collaborative filtering in recommender systems (Sarwar et al., 2001), where scalable approaches are studied extensively. We start by setting up the problem formulation of predicting example forgetting, and evaluate the performance of prediction of different approaches. We present reduced forgetting by utilizing the prediction outcomes during fine-tuning.
@@ 4.1 Training and Evaluation of Forgetting Prediction
Figure 5: The training and testing setup of predicting example forgetting with association matrix completion.
To apply matrix completion for predicting forgetting, a few entries $z_{ij}$ should be known when a new fine-tuning task $T_{i}\in\mathcal{T}_{\textrm{test}}$ (row $i$) is introduced. We therefore assume access to the ground truth forgetting $z_{ij}$ of a tiny random set $\mathcal{S}$ ($|\mathcal{S}|=30$) of upstream examples for $T_{i}\in\mathcal{T}_{\textrm{test}}$, noted as seed forgetting $z_{i}^{\mathcal{S}}=\{z_{ij}|x_{j}\in\mathcal{S}\}$. Obtaining seed forgetting typically takes only a few seconds by running inference with the model fine-tuned on $T_{i}$ over $\mathcal{S}$; we then predict forgetting of the rest $10k-100k$ upstream examples. Figure 5 illustrates an example of the train-test partition, seed forgetting, and the forgetting to be predicted. We use Root Mean Squared Error (RMSE) over the $\mathcal{T}_{\textrm{test}}$ as the metrics of predicting example forgetting.
Matrix completion approaches. We run matrix completion algorithms including additive linear, SVD  [^2] We improved the results of SVD in Table 2 after more hyperparameter tuning in this version., and k-nearest neighbors (KNN) models. The additive linear and the SVD models are introduced earlier in Sec. 3. Given the seed forgetting $z_{i}^{\mathcal{S}}$ of a task $T_{i}\in\mathcal{T}_{\textrm{test}}$, KNN finds tasks from $\mathcal{T}_{\textrm{train}}$ that have similar patterns of forgetting over the seed upstream examples $\mathcal{S}$. KNN computes an average of forgetting of top-k similar tasks from $\mathcal{T}_{\textrm{train}}$ weighted by their similarity as the prediction of forgetting caused by $T_{i}\in\mathcal{T}_{\textrm{test}}$ on the upstream examples $x_{1..M}$.
@@
We leave the implementation details of matrix completion approaches and the learned similarity approach in Appendix B.
Table 2: RMSE of predicting example forgetting over a held-out set of upstream examples after fine-tuning LMs on unseen new tasks. We report averaged performance over different seed sets ($\mathcal{S}$) of upstream examples with known ground truth forgetting beforehand.
In-Domain Out-of-Domain
OLMo-1B OLMo-7B MPT-7B OLMo-7B-Inst. OLMo-1B OLMo-7B MPT OLMo-7B-Inst.
Full FT Full FT LoRA Full FT Full FT LoRA Full FT Full FT LoRA Full FT Full FT LoRA
Additive 2.81 7.40 3.50 13.33 15.57 6.12 2.81 5.83 7.01 10.02 38.90 21.22
SVD 2.80 7.14 3.48 10.41 13.74 5.89 2.82 5.76 6.80 7.03 40.47 20.23
KNN 2.79 7.33 3.45 12.80 14.30 5.54 2.84 5.83 6.83 7.71 38.77 20.82
Similarity 3.84 9.29 5.45 14.00 16.23 6.19 3.93 7.64 8.52 10.97 42.38 23.47
4.2 Mitigating Forgetting with Predicted Forgetting
As we have discussed in Sec. 4.1, predicting forgetting with matrix completion requires seed forgetting $z^{S}$ to be evaluated. We consider an offline and an online variant of the approach. The offline variant performs a replay-free run of fine-tuning on the task $T_{i}$, after which the seed forgetting will be evaluated. We then perform another run of fine-tuning while replaying examples with the predicted forgetting. This creates computational overhead equivalent to one extra run of fine-tuning, but is still efficient when the training set of fine-tuning is considerably smaller than the upstream data. The online variant instead replays random examples for first 10% of fine-tuning steps, after which it evaluates seed forgetting and determine examples to be replayed in the rest of 90% steps. Compared to the offline variant, this mitigates the extra overhead of fine-tuning by trading off the prediction accuracy of forgetting.
Baselines of mitigating forgetting. We compare with diverse strategies of selecting upstream examples for sparse replay. We primarily examine whether weighted sampling with predicted forgetting statistically significantly improves over random sampling of upstream examples (Random). We also compare with Maximally Interfered Retrieval (MIR) (Aljundi et al., 2019a), a selection strategy sharing the similar notion of importance that forgotten examples should be selected for replay. The approach performs bi-level sampling by selecting the most forgotten examples from a small random subset of upstream data (set as $|\mathcal{S}|=30$). In addition, we apply strategies that consider different definitions of upstream example importance. We examine an approach based on perplexity thresholds (PPL) (Marion et al., 2023), which samples upstream data of which the perplexity is around the median of the distribution. For OLMo-1B, we also sample replayed examples proportional to the gradient inner products (Grad-Prod) we evaluated in Sec. 3.3, in a similar vein to coreset selection approaches that utilize gradient information (Park et al., 2023; Xia et al., 2024). As a reference, we also experiment with upweighting upstream examples with ground truth forgetting $z_{ij}$, which, however, can face computational efficiency issues in practice.
4.3 Results of Predicting and Mitigating Forgetting
Results of predicting example forgetting. Table 2 summarizes the error of predicting example forgetting over tasks from the in-domain and out-of-domain test splits. We see matrix completion approaches consistently outperform the learned similarity model in the prior work. Among the three matrix completion approaches, we notice that SVD models in general achieve the lowest prediction error. Besides, KNN in general outperforms additive linear and learned similarity models while being highly computationally efficient. Given this finding, we primarily leverage predictions given by the KNN and SVD models to reduce forgetting during fine-tuning.
@@ 4.3 Results of Predicting and Mitigating Forgetting
Figure 6: Log perplexity over upstream data as replay upstream examples selected by different approaches after fine-tuning OLMo or OLMo-Instruct models over in-domain or out-of-domain test tasks. The solid horizontal lines indicate the log perplexity before fine-tuning (i.e., no forgetting). The dash lines show the log perplexity achieved by upweighting upstream examples according to their ground truth forgetting. * and ** indicate statistical significance of improvement ($p<0.05$ or $p<0.005$) compared to replaying random examples in paired $t$-tests over all fine-tuning tasks.
@@
Mitigating forgetting with the predicted forgetting. We leverage the online or offline predicted forgetting by the KNN and SVD models to reweight examples during replay following our procedure in Sec. 4.2. Figure 6 summarizes log perplexity after fine-tuning over the held-out (never replayed) upstream data as we apply different upstream example selection approaches. We visualize the relative comparisons of these approaches to no forgetting of upstream examples (log perplexity before fine-tuning). We notice that example selection based on gradient inner products (Grad) or perplexity threshold (PPL), mainly applied for identifying important training data for a task in prior works, does not show improvement in mitigating forgetting compared to replaying random examples. This implies that the notions of example importance in these works are different from how easily the examples are forgotten. We also notice that MIR does not improve over random sampling in our setup, likely because of the small size of the retrieval candidates relative to the upstream examples. Upweighting examples with ground truth forgetting (GT) consistently reduces forgetting compared to random examples. By utilizing predicted forgetting by offline SVD and KNN we statistically significantly reduce forgetting compared to random examples in 4 out of 5 tested setups. The nuanced differences between SVD and KNN in forgetting mitigation align with their in-domain prediction accuracy of forgetting in Table 2, where SVD outperforms KNN by most on OLMo-7B models. Utilizing online-predicted forgetting also statistically significantly improves over replaying random examples in 3 of the setups. The gaps between online and offline variants are closer on 7B models than 1B models.
@@
Effects on downstream task performance measured with task-specific metrics. We evaluate fine-tuned OLMo-7B or OLMo-7B-Instruct on unseen LLM leaderboard tasks and present the results in Appendix C. We notice the performance stays stable or improves on some downstream tasks while degrades on others, indicating forgetting. Although we observe slightly improved performance of Offline-KNN to random or no replay on most forgotten tasks (e.g., Sciq on OLMo-7B and IFEval on OLMo-7B-Instruct), we do not see statistical significance. We leave effective algorithms to mitigate downstream task forgetting with predicted forgetting as future work.
@@
Table 3: Computational cost of replay-based approaches as a summation of fine-tuning costs $FT(\cdot)$, inference costs over upstream examples $EV(\cdot)$, and matrix completion costs $MC$.
@@
Offline SVD, KNN $2FT(Y)+EV(S)+MC$
Online SVD, KNN $FT(Y)+EV(S)+MC$
MIR $2FT(Y)+Y\cdot EV(S)$
@@
Computational Efficiency. Table 3 summarizes the computation cost of the approaches as a function $FT(\cdot)$ of fine-tuning steps, a function $EV(\cdot)$ of upstream examples whose perplexity is evaluated, and the cost of matrix completion (MC) that is much smaller than LLM inference or training. We note the total number of upstream examples as $N$, the size of seed examples as $S$, and the number of fine-tuning steps as $Y$. As $S$ is much smaller than $M$, the majority of computational costs arise from fine-tuning $FT(Y)$ and $EV(N)$. Replaying with ground truth forgetting is the most costly, as it introduces an additional run of fine-tuning (after which forgetting will be evaluated) and inference over potentially very large-scale upstream data. As offline prediction of forgetting only mitigates the need for inference, the approach saves computations when the cost of fine-tuning $FT(Y)$ is notably smaller than $EV(N)$, i.e., over small fine-tuning datasets and massive upstream data. Online prediction of forgetting is always efficient, requiring only one run of fine-tuning and without the need to evaluate over massive upstream data.

### v4

@@ 3.3 Interpreting Example Associations
Figure 5: Patterns in associations captured by $k$-th component in the matrix factorization of the example associations $Z$, exemplified with OLMo-7B-Instruct (full-parameter tuning). The highlighted regions display patterns that TruthfulQA tasks cause more forgetting ($k=2$) and MMLU causes more forgetting on certain tasks in FLAN v2 and Open-Orca ($k=3$).
Patterns of forgetting from matrix factorizations. The matrix factorization of $Z$ yields interpretable patterns of forgetting in each of its component $\bm{\alpha}_{k}\bm{\beta}_{k}^{T}$. Figure 5 provides an example of patterns captured by the $k$-th component in OLMo-7B-Instruct experiments. Similarly, we analyze forgetting on OLMo-1B and 7B and summarize interpretations of each component in Appendix E. There are interesting yet semantically intriguing associations. For example, on OLMo-7B, Stackoverflow documents are less forgotten when learning summarization tasks, while more forgotten when learning certain paraphrasing tasks.
@@
Correlations between example associations and similarity measures. We examine whether the associations between learned tasks and forgotten examples are interpretable from the similarity between the learned tasks and upstream examples. We consider (1) heuristic similarity measures, such as token or representational similarity, and (2) first-order approximations, such as inner products of gradients and inner products between weight updates and gradients (Lee et al., 2019; Doan et al., 2020). We detail each similarity measure in Appendix F. We then evaluate correlations between the actual forgetting $z_{ij}$ and the various similarity measures on OLMo-1B and summarize the results in Table 1.
Forgetting correlates poorly with similarity measures of learned tasks and upstream examples. From Table 1, we notice that none of the similarity measures correlates strongly with the actual forgetting, with a correlation $|\rho|<0.1$. These results imply that although simple statistical patterns are dominant in the example forgetting, such associations are not well-interpreted with common similarity measures of learned tasks and forgotten examples. Therefore, we hypothesize that leveraging the statistics of forgetting allows better prediction of forgetting than the contents of the tasks and examples.
@@
Table 1: Correlations between various measures of similarity and upstream example forgetting.
@@
[…]
@@ 4 Predicting Example Forgetting with Association Matrix Completion
We utilize our findings in Sec. 3 to predict example forgetting as the model learns a new task, a problem also studied in prior works (Jin & Ren, 2024). Predicting most forgotten examples allows better spot of the behavior changes of the models, enabling efficient and targeted approaches to mitigate forgetting, e.g., by replaying these examples. Although the ground truth forgetting can be directly obtained by running inference with the fine-tuned model over the upstream data, this requires extensive computational resources; in constrast, once a prediction model is trained, forgetting caused by unseen tasks can be predicted efficiently.
Following analysis in Sec. 3, we formulate prediction of example forgetting as a matrix completion problem over the empirical associations $Z$. We start by setting up the problem formulation of predicting example forgetting, and evaluate the performance of different matrix completion algorithms. We then demonstrate the practical benefit of predicting example forgetting by utilizing the prediction outcomes to mitigate forgetting during fine-tuning.
@@ 4.1 Training and Evaluation of Forgetting Prediction
Figure 6: The training and testing setup of predicting example forgetting with association matrix completion, and their integration into example replay methods to mitigate forgetting.
Table 2: RMSE ($\downarrow$) or F1($\uparrow$) of predicting example forgetting over a held-out set of upstream examples after fine-tuning LMs on unseen new tasks. We report average performance over different seed sets ($\mathcal{S}$) of upstream examples with known ground truth forgetting beforehand.
In-Domain Out-of-Domain
Model OLMo-1B OLMo-7B
OLMo-7B
LoRA
MPT
OLMo-7B
Instruct
OLMo-1B OLMo-7B
OLMo-7B
LoRA
MPT
OLMo-7B
Instruct
Metrics RMSE RMSE RMSE RMSE RMSE F1 RMSE RMSE RMSE RMSE RMSE F1
Additive 2.81 7.40 3.50 13.33 15.57 54.31 2.81 5.83 7.01 10.02 38.90 43.57
KNN 2.79 7.33 3.45 12.80 14.30 55.68 2.84 5.83 6.83 7.71 38.77 44.11
MF 2.80 7.14 3.48 10.41 13.74 57.35 2.82 5.76 6.80 7.03 40.47 42.91
Similarity 3.84 9.29 5.45 14.00 16.23 55.41 3.93 7.64 8.52 10.97 42.37 42.45
To apply matrix completion for predicting forgetting, a few entries $z_{ij}$ should be known when a new fine-tuning task $T_{i}\in\mathcal{T}_{\textrm{test}}$ (row $i$) is introduced. We therefore assume access to the ground truth forgetting $z_{ij}$ of a tiny random set $\mathcal{S}$ ($|\mathcal{S}|=30$) of upstream examples for $T_{i}\in\mathcal{T}_{\textrm{test}}$, noted as seed forgetting $z_{i}^{\mathcal{S}}=\{z_{ij}|x_{j}\in\mathcal{S}\}$. Obtaining seed forgetting typically takes only a few seconds by running inference with the model fine-tuned on $T_{i}$ (or fine-tuned for a few steps on $T_{i}$, which we evaluate separately) over $\mathcal{S}$. We then predict forgetting of the rest $10k-100k$ upstream examples. Figure 6 illustrates an example of the train-test partition, seed forgetting, and the forgetting to be predicted. We use Root Mean Squared Error (RMSE) or F1 over the $\mathcal{T}_{\textrm{test}}$ as the metrics of predicting log-perplexity increase or exact match drop.
Matrix completion approaches. We run matrix completion algorithms including additive linear models, matrix factorization (MF), and k-nearest neighbors (KNN) models. The additive linear model approximate forgetting as additive effects of learned tasks $\alpha^{(i)}$ and upstream examples $\beta^{(j)}$ ($\alpha^{(i)}+\beta^{(j)}$). The MF models are introduced earlier in Sec. 3. Given the seed forgetting $z_{i}^{\mathcal{S}}$ of a task $T_{i}\in\mathcal{T}_{\textrm{test}}$, KNN finds tasks from $\mathcal{T}_{\textrm{train}}$ that have similar patterns of forgetting over the seed upstream examples $\mathcal{S}$. KNN computes an average of forgetting of top-k similar tasks from $\mathcal{T}_{\textrm{train}}$ weighted by their similarity as the prediction of forgetting caused by $T_{i}\in\mathcal{T}_{\textrm{test}}$ on the upstream examples $x_{1..M}$.
@@
We leave the implementation details of matrix completion approaches and the learned similarity-based prediction approach in Appendix B.
@@
Figure 7: Log perplexity over upstream data with different strategies of selecting examples to be replayed. The solid horizontal lines indicate the log perplexity before fine-tuning (i.e., no forgetting). The dash lines show the log perplexity achieved by upweighting upstream examples according to their ground truth forgetting. * and ** indicate statistical significance of improvement ($p<0.05$ or $p<0.005$) compared to replaying random examples in paired $t$-tests over all fine-tuning tasks.
4.2 Mitigating Forgetting with Predicted Forgetting
As we have discussed in Sec. 4.1, predicting forgetting with matrix completion requires seed forgetting $z^{S}$ to be evaluated. We consider an offline and an online variant of the approach. The offline variant performs a replay-free run of fine-tuning on the task $T_{i}$, after which the seed forgetting will be evaluated. We then perform another run of fine-tuning while replaying examples with the predicted forgetting. This creates computational overhead equivalent to one extra run of fine-tuning, but is still efficient when the training set of fine-tuning is considerably smaller than the upstream data. The online variant instead replays random examples for first 10% of fine-tuning steps, after which it evaluates seed forgetting and determine examples to be replayed in the rest of 90% steps. Compared to the offline variant, this mitigates the extra overhead of fine-tuning by trading off the prediction accuracy of forgetting. We illustrate the two variants in Figure 6.
Baselines of mitigating forgetting. We compare with diverse strategies of selecting upstream examples for sparse replay. We primarily examine whether weighted sampling with predicted forgetting statistically significantly improves over random sampling of upstream examples (Rand). We also compare with a variant of Maximally Interfered Retrieval (MIR-T) (Aljundi et al., 2019a), a selection strategy sharing the similar notion of importance that forgotten examples should be selected for replay. The approach performs bi-level sampling by selecting the most forgotten examples from a small random subset of upstream data. We extend the approach to select forgotten examples after a full training run on a task, instead of single steps, which achieves better performance. In addition, we apply strategies that consider different definitions of upstream example importance. We examine an approach based on perplexity thresholds (PPL) (Marion et al., 2023), which samples upstream data of which the perplexity is around the median of the distribution. For OLMo-1B, we also sample replayed examples proportional to the gradient inner products (Grad-Prod) (whose correlation with forgetting is evaluated in Table 1), a representative coreset selection approach that utilizes gradient information (Park et al., 2023; Xia et al., 2024). As a reference, we also experiment with upweighting upstream examples with ground truth forgetting $z_{ij}$, which, however, is highly inefficient and repetitive to obtain in practice.
4.3 Results of Predicting and Mitigating Forgetting
Results of predicting example forgetting. Table 2 summarizes the error of predicting example forgetting over tasks from the in-domain and out-of-domain test splits. We see matrix completion approaches consistently outperform the learned similarity model in the prior work. Among the three matrix completion approaches, we notice that MF models in general achieve the lowest prediction error. Besides, KNN in general outperforms additive linear and learned similarity models while being highly computationally efficient.
@@ 4.3 Results of Predicting and Mitigating Forgetting
Mitigating forgetting with the predicted forgetting. We leverage the online or offline predicted forgetting by the matrix completion algorithms to reweigh examples during replay following our procedure in Sec. 4.2. Figure 7 summarizes log perplexity after fine-tuning over the held-out (never replayed) upstream data as we apply different upstream example selection approaches. We notice that example selection based on gradient inner products (Grad) or perplexity threshold (PPL), mainly applied for identifying important training data for a task in prior works, does not show improvement in mitigating forgetting compared to replaying random examples. This implies that the notions of example importance in these works are different from how easily the examples are forgotten. We also notice that MIR-T does not improve over random sampling in our setup, likely because of the small size of the retrieval candidates relative to the upstream examples. Upweighting examples with ground truth forgetting (GT) consistently reduces forgetting compared to random examples. By utilizing predicted forgetting by offline additive linear, KNN, and MF models, we statistically significantly reduce forgetting compared to random examples in 4 out of 5 tested setups. MF models achieves statistical significance in most the scenarios, which aligns with its top average prediction performance of example forgetting in Table 2. Utilizing online-predicted forgetting also statistically significantly improves over replaying random examples in 3 of the setups. The gaps between online and offline variants are closer on 7B models than 1B models.
Computational efficiency discussions. Table 3 summarizes the computation cost of the approaches as a function $FT(\cdot)$ of fine-tuning steps, a function $EV(\cdot)$ of upstream examples whose perplexity is evaluated, and the cost of matrix completion (MC) that is much smaller than LLM inference or training. We note the total number of upstream examples as $N$, the size of seed examples as $S$, and the number of fine-tuning steps as $Y$. As $S$ is much smaller than $M$, the majority of computational costs arise from the fine-tuning $FT(Y)$ and the upstream data evaluation $EV(N)$. Replaying with ground truth forgetting is the most costly, as it introduces an additional run of fine-tuning (after which forgetting will be evaluated) and inference over potentially very large-scale upstream data. The offline prediction and replay approach saves computations in the scenarios of small fine-tuning datasets and massive upstream data, which is often true in practice. Online prediction of forgetting does not incur extra cost of fine-tuning or upstream data inference, thereby always being efficient.
@@
Effects on downstream task performance measured with task-specific metrics. We evaluate fine-tuned OLMo-7B or OLMo-7B-Instruct on unseen LLM leaderboard tasks and present the results in Appendix D. Although we observe slightly improved performance of replaying predicted examples to random or no replay on most forgotten tasks (e.g., Sciq on OLMo-7B and IFEval on OLMo-7B-Instruct), we do not see statistical significance. We leave more effective algorithms to mitigate downstream task forgetting with predicted forgetting as future work, e.g., by integrating predicted forgetting into alternative algorithms that leverage past examples (Aljundi et al., 2017; Buzzega et al., 2020)
@@
Table 3: Computational cost of replay-based approaches as a summation of fine-tuning costs $FT(\cdot)$, inference costs over upstream examples $EV(\cdot)$, and matrix completion costs $MC$. Costs that are minor are displayed in smaller fonts.
@@
Offline MF $2FT(Y)+\scriptstyle{EV(S)}+\scriptstyle{MC}$
Online MF $FT(Y)+\scriptstyle{EV(S)}+\scriptstyle{MC}$
MIR-T $2FT(Y)+Y\cdot\scriptstyle{EV(S)}$